200 mL of 0.0010 M Sr(NO3)2 (aq) are mixed with 800 mL of 0.0050M NaF (aq) to make 1.0 L of solution. The Ksp value of SrF2 is 4.3 x 10-9. Which is correct? I know the answer, I just don't know how they got to it.

a. SrF2 will not precipitate.

b. SrF2 will precipitate. NaF is the limiting reactant.

c. SrF2 will precipitate. NaNO3 is the limiting reactant.

d. SrF2 will precipitate. Sr(NO3)2 is the limiting reactant.

e. SrF2 will precipitate. SrF2 is the limiting reactant.
Answer #1

200 mL of 0.0010 M Sr(NO3)2 (aq) = 0.200 L * 0.0010 mole / L = 2.0 * 10^-4 mole.

800 mL of 0.0050 M NaF (aq) = 0.800 L * 0.0050 mole / L = 4.0 * 10^-3 mole,

Sr^2+ + 2 F- ...........> SrF2

total volume = (200 +800) = 1000 ml = 1.00 L

thus

[Sr^2+] = 2.0 * 10^-4 mole / 1.0 L = 2.0 * 10^-4 M

and

[F-] = 4.0 * 10^-3 mole / 1.00 L = 4.0 * 10^-3 L

Q = [Sr^2+][F-]^2 = 2.0 * 10^-4 * (4.0 * 10^-3)^2 = 3.2 * 10^-9

we have

Ksp value of SrF2 is 4.3 x 10-9

as Ksp > Q

and

hence SrF2 will not precipitate.

a. SrF2 will not precipitate. is the answer.